Adv. Math. Econ. 15, 89–122 (2011)

On the perception and representation of economic quantity in the history of economic analysis in view of the Debreu conjecture ^{1}

Akira Yamazaki ^{2}

Professor, School and Graduate School of Economics, Meisei University, 2-1-1 Hodokubo, Hino, Tokyo 191-8506, Japan Emeritus Professor, Graduate School of Economics, Hitotsubashi University, 2-1, Naka, Kunitachi, Tokyo 186-8601, Japan (e-mail: yamazaki@econ.hit-u.ac.jp)

Received: October 19, 2010 Revised: November 3, 2010

JEL classification: B13, B23, B3

Mathematical Subject Classification (2010): 62P20, 91B02, 91B42, 91B50

Abstract. In this paper, the perception and representation of economic quantity found in the works of Augustin Cournot, Leon Walras, Vilfreto Pareto, and Alfred Marshall will be discussed. An interpretation of the perception and representation of economic quantities and economic variables, specifically relating to the concept of demand, in the fundamental theoretical framework of general equilibrium theory will be provided, particularly from the vantage point of the Debreu conjecture.

Key words: Debreu conjecture, economic quantity, demand, general equilibrium theory, A. Cournot, L. Walras, V. Pareto, A.Marshall

1This paper is the English version of an article written originally in Japanese. The Japanese version was published in both the Mita Journal of Economics 103(1), 2010, and in Epimetheus in Economics, edited by T. Maruyama (Chisen Shokan: Tokyo, 2010). The author is heavily indebted to Professor Philip Culbertson for suggesting many elegant improvements to the manuscript in English.

2I acknowledge the support for this research through a JSPS Grant in Aid for Scientific Research, no. 22530188.